OpenStudy (anonymous):

(Algebra 1 Question) A test has twenty questions worth 100 points. The test consists of x true-false questions worth 4 points each and y multiple choice questions worth 8 points each. How many of each type of question are on the test?

OpenStudy (retireed):

You have two unknowns x true false and Y multiple choice So you need two equations to solve. Can you give me one equation?

OpenStudy (anonymous):

I'm not sure I understand, but 4x+8y=100?

OpenStudy (retireed):

Very good! The other is much easier to figure out.

OpenStudy (retireed):

It involves the total number of questions.

OpenStudy (anonymous):

I really dont think there is another one...

OpenStudy (retireed):

Yes there is, think of the total number of questions.

OpenStudy (anonymous):

x+y=20?

OpenStudy (anonymous):

All i need here is the second equation because I can graph the rest

OpenStudy (anonymous):

I dont know how to find it

OpenStudy (anonymous):

4x+8y=20?

OpenStudy (anonymous):

actually that equation would make sense...

OpenStudy (retireed):

Now you have two equations which you can solve for x or y.

OpenStudy (anonymous):

okay, thanks. I dont need to solve anything, they gave me a graph xD

OpenStudy (anonymous):



OpenStudy (retireed):

you found them both 4x+8y=100 and x+ y = 20

OpenStudy (retireed):

You're welcome.